OpenStudy (anonymous):

How do i turn y = ax2 + bx + c to (h, k) . and how do i turn (h, k) to ax2 + bx + c

OpenStudy (anonymous):

help

OpenStudy (amistre64):

what does (h,k) signify?

OpenStudy (anonymous):

the vertex

OpenStudy (amistre64):

oh, then it makes sense now

OpenStudy (amistre64):

the vertex is olong the axis of symmetry defined to be: x = -b/(2a)

OpenStudy (amistre64):

or are you trying to convert it to a(x-h)^2 + k algebraically?

OpenStudy (anonymous):

algebraically

OpenStudy (campbell_st):

(h, k) is the vertex, I'd expect... so group the terms in x and then complete the square here is an example \[y = x^2 + 6x - 7\] complete the square in x \[y = (x^2 + 6x + 9) - 7 - 9\] adding 9 makes a perfect square in x, but to keep the equation in balance 9 also needs to be subtracted. so it becomes \[y=(x + 3)^2 - 16\] with vertex at (-3, -16) so h = -3 and k = -16 hope it helps

OpenStudy (amistre64):

simplest thing to do is expand it out and compare parts a(x^2 -2xh + h^2) + k ax^2 -2axh + ah^2 + k -2ah = b ah^2+k = c

OpenStudy (anonymous):

where the 9 come from

OpenStudy (amistre64):

by comparison h = -b/(2a) k = c - ah^2 = c -ab^2/(4a^2) = c -b^2/(4a) = (4ac -b^2)/(4a)

OpenStudy (anonymous):

still confused, anyone got a video or something

OpenStudy (amistre64):

you tube completing the square if thats what helps you learn the best

OpenStudy (anonymous):

jow did u get from (x2+6x+9) to (x+3)2

OpenStudy (anonymous):

how